OpenStudy (anonymous):

Find equation of parabola if vertex at (−4, 2), axis parallel to the x-axis and the line x = 6− 4y is a tangent.

OpenStudy (anonymous):

@jim_thompson5910 Help again please?

OpenStudy (anonymous):

(y-k)^2 = 4a(x-h) (y-2)^2 = 4a(x+4)

jimthompson5910 (jim_thompson5910):

so far so good, just have to figure out what to do with that tangent one sec

OpenStudy (anonymous):

would be sub in (−4, 2) into tangent? like to get (-2, 5/2)?

jimthompson5910 (jim_thompson5910):

that's the thing, the tangent isn't at (-4,2)

OpenStudy (anonymous):

well then i have no idea what to do with the tangent :(

jimthompson5910 (jim_thompson5910):

same here, but trying something out

OpenStudy (anonymous):

well, im gonna take a shower now, brb

OpenStudy (anonymous):

we got the same name O_O

jimthompson5910 (jim_thompson5910):

does it say where the tangent line is tangent at? like where the tangent point is?

OpenStudy (anonymous):

you would find it by equating x = 6− 4y and whatever the formula is for the parabola

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

but does that guarantee that x = 6− 4y is a tangent?

OpenStudy (phi):

If you know calculus, you can solve for dy/dx of the parabola and set that equal to the slope of the line. you can then find x and y as a function of a, substitute those values into the parabola and solve for a.

OpenStudy (anonymous):

how would you do that? we dont even know the parabola.

OpenStudy (phi):

do you know calculus ?

OpenStudy (anonymous):

yea

OpenStudy (phi):

take the derivative with respect to x of (y-2)^2 = 4a(x+4) 2(y-2) dy/dx = 4a dy/dx = 2a/(y-2) find the slope of the line x = 6− 4y is: 4y= -x + 6, y = -1/4 x +6/4 m= -1/4 -1/4 = 2a/(y-2) solve for y use x = 6− 4y to solve for x in terms of a use those in (y-2)^2 = 4a(x+4) to find a

OpenStudy (phi):

Here's a graph

OpenStudy (anonymous):

so i got x= 6-4y y= -8a + 2

OpenStudy (phi):

x= 6-4y y= -8a + 2 so x= 6 -4(-8a+2) = 6 +32a-8 = 32a-2 use those in (y-2)^2 = 4a(x+4)

OpenStudy (anonymous):

how can we use those in (y-2)^2 = 4a(x+4)??

OpenStudy (phi):

substitute y= -8a+2 and x= 32a-2 into (y-2)^2 = 4a(x+4)

OpenStudy (anonymous):

ahh ok :)

OpenStudy (tkhunny):

Vertex: (-4,2) Axis of Symmetry parallel to the x-axis Tangent to x = 6-4y This restricts the solution to one type. Opens in the negative x-direction. Focus: (-4-a,2) for some a > 0 Directrix: x = -4 + a Parabolas is all ordered pairs (x,y) such that the distance to the focus is the distance to the directrix. Distance to Focus: \(\sqrt{(x+4+a)^{2}+(y-2)^{2}}\) Distance to Directrix: x + 4 - a Equating these and solving gives: \(-4a(x+4) = (y-2)^{2}\) Whix is good for \(x \le -4\) Now for the line: x + 4y - 6 = 0 Distance to Directrix: x + 4 - a (Not very helpful, yet.) Distance to Focus: \(\dfrac{|1(-4-a) + 4(2) - 6|}{\sqrt{1^{2}+4^{2}}} = \dfrac{|a+2|}{\sqrt{17}} = \dfrac{a+2}{\sqrt{17}}\), since a > 0, we can remove the absolute values. The we know: \(x + 4 - a = \dfrac{a+2}{\sqrt{17}}\) This is the third required constraint.

OpenStudy (tkhunny):

Note: This is why we invented the calculus!